# AMMeBa: A Large-Scale Survey and Dataset of Media-Based Misinformation In-The-Wild

### v1

@@ Abstract
The prevalence and harms of online misinformation is a perennial concern for internet platforms, institutions and society at large. Over time, information shared online has become more media-heavy and misinformation has readily adapted to these new modalities. The rise of generative AI-based tools, which provide widely-accessible methods for synthesizing realistic audio, images, video and human-like text, have amplified these concerns. Despite intense interest on the part of the public and significant press coverage, quantitative information on the prevalence and modality of media-based misinformation remains scarce. Here, we present the results of a two-year study using human raters to annotate online media-based misinformation, mostly focusing on images, based on claims assessed in a large sample of publicly-accessible fact checks with the ClaimReview markup. We present an image typology, designed to capture aspects of the image and manipulation relevant to the image’s role in the misinformation claim. We visualize the distribution of these types over time. We show the the rise of generative AI-based content in misinformation claims, and that it’s commonality is a relatively recent phenomenon, occurring significantly after heavy press coverage. We also show “simple” methods dominated historically, particularly context manipulations, and continued to hold a majority as of the end of data collection in November 2023. The dataset, Annotated Misinformation, Media-Based (AMMeBa), is publicly-available, and we hope that these data will serve as both a means of evaluating mitigation methods in a realistic setting and as a first-of-its-kind census of the types and modalities of online misinformation.
Figure 1: Examples media occurring alongside fact-checked misinformation claims. In this report, we introduce a typology to capture the enormous variation in media-based (particularly image-based) misinformation seen in-the-wild and categorize a very large sample of misinformation claims with it.
@@ 1 Summary of Findings
  • Image manipulations in the annotated claims were generally simple and do not require technological sophistication to produce; the most common type are context manipulations, which use frequently authentic images alongside a false claim about what they depict.
@@ 2 Introduction
Figure 2: Media manipulations have a long history. Top Left: A comparison of an image of Joseph Stalin, originally taken in 1937, where where an associate, Nikolai Yezhov, is present along with a later version where he has been manually removed from the official image with airbrushing, following his fall from favor. Top Middle A still from the 1967 “Patterson-Gimlin film,” alleged to show the cryptid Bigfoot, taken along the Klamath River in California. While it’s authenticity is disputed to this day, experts regard the footage as a hoax. Top Right A 1994 Time Magazine cover featuring a mugshot of O.J. Simpson that had been artificially darkened to appear more sinister, compare the same photograph on the cover of Newsweek. Bottom Left The “Tourist Guy” image, originally taken in 1997. The creator of the image confessed to manipulating the image to add the airplane “as a joke.” Bottom Middle A 2008 image of Kim Jong Il that a BBC analysis concluded was fake. Bottom Right A 2024 AI-generated image, dubbed “Shrimp Jesus,” that went viral on social media.
@@
The bulk of the annotations center in particular on misinformation claims that involve images. The characterization, and the associated typology, are oriented around the mechanism by which the images are used to bolster the misinformation claim: the inclusion of text, the type of manipulation, etc. As such, annotations address on the types of manipulations performed and how they are realized, rather than the subject matter of the claims themselves, to ensure the information is instrumentally useful to those working in developing tools and mitigation methods that rely on the properties of the media or its context.
@@ 3 Related Work
The rise of generative AI methods has motivated a number of image-based datasets, which consist of AI-generated images, sometimes with the generator prompt used to obtain the image. Such datasets are very frequently not only consist of synthetic data but are also produced in a synthetic way, by direct sampling large numbers of images, for instance GenImage [39]. Others, however, do capture in-the-wild distributions. Two such datasets collect images and prompts from channels on the chat app Discord [^7] https://discord.com/ that are specific to popular image generators Midjourney [^8] https://www.midjourney.com/home [1] and Stable Diffusion [^9] https://stability.ai/ (DiffusionDB) [35]. While sampled from in-the-wild sources, both restrict their attention to a single synthesizer. TWIGMA [8], whose samples are inferred to be synthetic based on metadata, uses Twitter [^10] https://twitter.com/ (now “X”) as a source and thus aims to be more varied and comprehensive and should emulate the distribution of generators and synthesis methods used online.
@@
Misinformation often depends on the interaction between text and media. COSMOS [4] provides image and text pairs and synthetically create “context manipulations” (5.4.2). McCrae et al. [20] construct a similar dataset but using videos instead of images, collected from Facebook [^12] https://www.facebook.com/. McCrae et al. and use real (video, caption) pairs as authentic examples, but like COSMOS construct synthetic manipulations by permuting those pairs. These satisfy the technical definitions of the context manipulations (which are often deployed in misinformation, see Sec. 7.5) they seek to model but they lack the intent and bespoke curation seen in authentic samples.
@@
An alternative, less content-specific approach is to restrict analyses to a specific platform. Grinberg et al. [12] study the spread of misinformation on Twitter leading up to the 2016 United States presidential election, focusing primarily on the frequency of exposure to fake news by Twitter users, particularly as a function of political affinity, demographics, and consumption of political content overall. Addressing visual misinformation specifically, Yang et al. [38] manually annotate a random sample of political images posted to Facebook, and find a large contingent contain misinformation elements. Yang et al. additionally report “four main types” of misleading image posts (doctored images, memes with misleading text, images associated with misleading posts, and text-only images). These categories map fairly neatly onto those outlined in this study. However, they do not provide further details beyond this enumeration.
@@ 5.1 Top Level Types
This is a partially subjective judgement; consensus was reached after several rounds of trial labelling and review.
@@ 5.3.2.1 Screenshots
Screenshots are a class of complex images that include GUI elements that conclusively show the image was captured by a user taking a screenshot of content on an electronic display. In order to qualify for this category, the images must contain GUI elements that are unambiguous and recognizable. While a screenshots are a type of complex image, they are broken out here because of their prevalence and the differences in the investigatory approach compared to general complex images.
@@ 5.4 Manipulation Types
  • Text-based images do not require the visual content of an image possesses any specific qualities. They are misinformation claims, as text, rendered in an image. The image may feature other content, but that content is not important for the claim being made. Images which feature no misinformation-relevant non-textual visual content at all are a common element of this category.
@@ 5.4.1 Content Manipulations
Content manipulations are defined here as taking place when (i) the content of the image itself has been altered or synthesized, (ii) it appears endogenous or “native” to the image itself (i.e., not clearly overlaid or added after-the-fact [^15] This is a generous criteria; even very crude attempts to make the manipulations appear present in the original are included here.) and (iii) the alteration made is relevant to the misinformation claim. Content manipulations have existed since the inception of photography, and are perhaps the canonical form of image-based misinformation in popular consciousness. They are created using a very wide variety of techniques, which are loosely situation onto a spectrum between Deepfakes (i.e., AI-generated) and Cheapfakes (“classical,” pre-AI methods) [26].
@@ 5.4.4 Fake Official Documents
This category is characterized by the presence of logos, letterhead, or other attributes that lend the false appearance of an official communication by a government, company, or other well-known organization or official body. They may also appear as articles from reputable news sources, either as physical news clippings or screenshots. Examples of fake official documents are given in Fig. 9. More detail on this category are available in the appendix (Sec. B.3.1).
@@ 6.2 Raters
A total of 83 raters participated in the study. Because of the complexity of the task, careful training and retention was prioritized over a large rater pool. Mean rater tenure was 147 days, performing an average of 5,971 annotation (counting annotations that were subsequently dropped or excluded from the analysis due to changes methodology). Two groups of raters were used simultaneously, one located in India and the other located in the United States. The groups were preferentially allocated fact checks sourced from their respective countries so that annotations would benefit from contextual knowledge; fact checks from neither country were allocated randomly.
@@ 7.1 Use of Media in Misinformation Claims
Fig. 11 plots the percentage of misinformation claims that rely on media using a 120-day sliding window based using an approximate publication date. While the percentage of media-based misinformation claims among all misinformation claims has been largely stable at least since 2019, such claims represent a large majority of around 80%, at least among those that are subject to a fact check. The cause of the decline between 2017 and 2019 is unclear; and may be due to changing prioritization in content addressed by fact checkers. The date of the fact check is used to reckon the date of the misinformation claim in this plot, and all subsequent plots containing data information. The fact check date is based on a combination of self-reported dates provided by fact checkers as part of the ClaimReview markup schema and the date when the ClaimReview-bearing fact check was encountered [^16] The dates provided in the dataset release are based purely on the self-reported dates from the ClaimReview schema..
@@
The secular increase in video-based misinformation is most apparent as a windowed percentage of total media-based misinformation addressed in fact checks over time. Viewed cumulatively in Fig. 13 the rise in video-based misinformation does not appear to occur alongside a decrease in other modalities. This suggests increases in fact checker capacity is preferentially allocated to video-based misinformation.
@@ 7.3 Manipulation Types
Our study provisions no measure of the comparative effectiveness or impact of these methods. Despite this, we do not regard the simplicity of producing a particular manipulation as a measure of its ineffective or benign nature; rather, the consistent popularity of context manipulation suggests their enduring effectiveness as a means of making a compelling false claim.
@@ 7.4 Content Manipulations
The prevalence of content manipulation sub-types (Sec. 5.4.1), as measured against the proportion of content manipulation types overall, is shown in Fig. 18 and has remained largely stable up until 2023. General (e.g., so-called “photoshops”) and text (applied specifically edits made to text that appears to occur on an object in the scene) manipulations are similarly common and make up a significant majority. Chyron manipulations are a relatively small proportion overall. AI-generated images made up a minute proportion of content manipulations overall until early last year. Starting shortly before 2023, generative AI-generated images begin to rapidly rise as a proportion of overall fact-checked image manipulations; such that the aggregate population of AI-generated images is now far greater than that of manipulated chyrons and nearly as common as text and general content manipulations. Note that the accuracy of the total count of fact-checked AI-content is subject in part to the completeness of their corresponding fact checks (Sec. 9.1): if the fact check does not indicate it’s AI, it will not be recorded as such by the rater.
@@
Figure 19: The “Hype Index” among content manipulations types in image-based misinformation over time. The “Hype Index” is computed as the mean number of fact checks per image over time, separated by content manipulation type. Inset is an AI-generated image of Pope Francis in a white jacket, which went viral in March of 2023 and was subject to numerous independent fact checks as well as being widely featured in the press in general. We interpret this the number of fact checks an image is subject to as an indicator of the fact checker community’s view of the importance of fact checking it. Under this interpretation, the massive spike in the index for AI-generated content indicates that fact checkers viewed communicating the nature of these images was of particular public interest (or at leas, of interest to the public) in early 2023. Plots are otherwise computed as in Fig. 11, although the minimum number of datapoints per interval required for inclusion has been changed from 100 to 10 to highlight the sharp spike in fact checks per AI image.
@@ 7.5 Context Manipulations
Images that are asserted by the fact check to be subject to context manipulations but which do not fall into the categories defined in Sec. 5.4.2 are not further characterized, and account for about 40% of all context manipulations. This reflects relatively stringent criteria for inclusion in a subcategory, and the breadth of what is considered context (more complete categorization of context manipulations is identified as an avenue for further research in Sec. 9.5). For example, a context manipulation can falsely state the identity of the originator of an image, but not of anything in the image, and so it would not fall under the identity manipulation subcategory [^17] Not capturing these additional subtypes of context manipulation is due mostly to the need to contain the complexity of the overall rater task.. These prevalence of these categories are shown in Fig. 20.
@@ 7.6 Image Text
Raters noted the presence of text, and did not discriminate between cases where the text occurs on an object in the scene or is digitally overlaid on top of the image. When text was both present and legible, or when a transcript was provided by the fact check (especially in the event the text was in a non-English language), raters were asked to assess whether or not the text was also relevant to the misinformation. Text relevance was determined using criteria similar to those used for the relevance of the image itself: some component of the misinformation claims depends on, or is supported by, the text content in the image. This dependence need not be particularly strong, and the category is used to separate cases where all text that occurs in an image is incidental to the misinformation claim its associated with.
@@
Figure 22: Self-contextualizing images. The prevalence of ‘self-contextualizing images,’ images subject to context manipulations where the text making articulating the false context is present in the image itself, among the total population of context manipulations. Plot created as in Fig. 11.
@@ 7.8 Fake Official Document
Fake official documents appear to be or to contain an official communication of some form from a well-known (or made to appear well-known) and reputable organization. The frequency of fake official documents is plotted in Fig. 24. Such images are of concern as they can disseminate false information during crises, masquerading as an official dispatch containing important information. Particularly salient examples occurred during the COVID-19 pandemic, where misinformation-bearing images would regularly occur appearing to be official information or instruction from a public health body. “Vaccine exception cards” and the like also fall into this category.
@@ 7.10 Effects of Global Events
The volume of fact checks correlates somewhat with global events. The period between the first recorded instance of COVID-19 in the United States on Jan 20, 2020 and the administration of the first of the vaccines on December 14, 2020 corresponds to the the most rapid increase in fact checks since the introduction of ClaimReview. To measure the effect on the distribution of different media types, we checked a battery of notable events against the observed trends. We find little evidence that the events themselves drive shifts in the composition of the misinformation-associated media population. Rather, trends appear to be driven by other phenomena, like the increasing popularity of video platforms or availability of generative-AI methods.
In some cases, changes are observed consistent with a priori expectations. For instance, the “Fake Official Documents” category experiences spikes in prevalence following the the deployment of COVID-19 tests and vaccinations. However, these observations are by no means conclusive.

### v2

@@ Abstract
The prevalence and harms of online misinformation is a perennial concern for internet platforms, institutions and society at large. Over time, information shared online has become more media-heavy and misinformation has readily adapted to these new modalities. The rise of generative AI-based tools, which provide widely-accessible methods for synthesizing realistic audio, images, video and human-like text, have amplified these concerns. Despite intense public interest and significant press coverage, quantitative information on the prevalence and modality of media-based misinformation remains scarce. Here, we present the results of a two-year study using human raters to annotate online media-based misinformation, mostly focusing on images, based on claims assessed in a large sample of publicly-accessible fact checks with the ClaimReview markup. We present an image typology, designed to capture aspects of the image and manipulation relevant to the image’s role in the misinformation claim. We visualize the distribution of these types over time. We show the rise of generative AI-based content in misinformation claims, and that its commonality is a relatively recent phenomenon, occurring significantly after heavy press coverage. We also show “simple” methods dominated historically, particularly context manipulations, and continued to hold a majority as of the end of data collection in November 2023. The dataset, Annotated Misinformation, Media-Based (AMMeBa), is publicly-available, and we hope that these data will serve as both a means of evaluating mitigation methods in a realistic setting and as a first-of-its-kind census of the types and modalities of online misinformation.
Figure 1: Examples of media occurring alongside fact-checked misinformation claims. In this report, we introduce a typology to capture the enormous variation in media-based (particularly image-based) misinformation seen in-the-wild and categorize a very large sample of misinformation claims with it.
@@ 1 Summary of Findings
  • Image manipulations in the annotated claims were generally simple and do not require technological sophistication to produce; the most common type are context manipulations, which use (frequently unmodified) images alongside a false claim about what they depict.
@@ 2 Introduction
Figure 2: Media manipulations have a long history. Top Left: A comparison of an image of Joseph Stalin, originally taken in 1937, where an associate, Nikolai Yezhov, is present along with a later version where he has been manually removed from the official image with airbrushing, following his fall from favor. Top Middle A still from the 1967 “Patterson-Gimlin film,” alleged to show the cryptid Bigfoot, taken along the Klamath River in California. While it’s authenticity is disputed to this day, experts regard the footage as a hoax. Top Right A 1994 Time Magazine cover featuring a mugshot of O.J. Simpson that had been artificially darkened to appear more sinister, compare the same photograph on the cover of Newsweek. Bottom Left The “Tourist Guy” image, originally taken in 1997. The creator of the image confessed to manipulating the image to add the airplane “as a joke.” Bottom Middle A 2008 image of Kim Jong Il that a BBC analysis concluded was fake. Bottom Right A 2024 AI-generated image, dubbed “Shrimp Jesus,” that went viral on social media.
@@
The bulk of the annotations center in particular on misinformation claims that involve images. The characterization, and the associated typology, are oriented around the mechanism by which the images are used to bolster the misinformation claim: the inclusion of text, the type of manipulation, etc. As such, annotations address the types of manipulations performed and how they are realized, rather than the subject matter of the claims themselves, to ensure the information is instrumentally useful to those working in developing tools and mitigation methods that rely on the properties of the media or its context.
@@ 3 Related Work
The rise of generative AI methods has motivated a number of image-based datasets, which consist of AI-generated images, sometimes with the generator prompt used to obtain the image. Such datasets very frequently not only consist of synthetic data but are also produced in a synthetic way, by direct sampling large numbers of images, for instance GenImage [39]. Others, however, do capture in-the-wild distributions. Two such datasets collect images and prompts from channels on the chat app Discord [^7] https://discord.com/ that are specific to popular image generators Midjourney [^8] https://www.midjourney.com/home [1] and Stable Diffusion [^9] https://stability.ai/ (DiffusionDB) [35]. While sampled from in-the-wild sources, both restrict their attention to a single synthesizer. TWIGMA [8], whose samples are inferred to be synthetic based on metadata, uses Twitter [^10] https://twitter.com/ (now “X”) as a source and thus aims to be more varied and comprehensive and should emulate the distribution of generators and synthesis methods used online.
@@
Misinformation often depends on the interaction between text and media. COSMOS [4] provides image and text pairs and synthetically create “context manipulations” (5.4.2). McCrae et al. [20] construct a similar dataset but using videos instead of images, collected from Facebook [^12] https://www.facebook.com/. McCrae et al. use real (video, caption) pairs as authentic examples, but like COSMOS construct synthetic manipulations by permuting those pairs. These satisfy the technical definitions of the context manipulations (which are often deployed in misinformation, see Sec. 7.5) they seek to model but they lack the intent and bespoke curation seen in authentic samples.
@@
An alternative, less content-specific approach is to restrict analyses to a specific platform. Grinberg et al. [12] study the spread of misinformation on Twitter leading up to the 2016 United States presidential election, focusing primarily on the frequency of exposure to fake news by Twitter users, particularly as a function of political affinity, demographics, and consumption of political content overall. Addressing visual misinformation specifically, Yang et al. [38] manually annotated a random sample of political images posted to Facebook, and find a large contingent contain misinformation elements. Yang et al. additionally report “four main types” of misleading image posts (doctored images, memes with misleading text, images associated with misleading posts, and text-only images). These categories map fairly neatly onto those outlined in this study. However, they do not provide further details beyond this enumeration.
@@ 5.1 Top Level Types
This is a partially subjective judgment; consensus was reached after several rounds of trial labeling and review.
@@ 5.3.2.1 Screenshots
Screenshots are a class of complex images that include GUI elements that conclusively show the image was captured by a user taking a screenshot of content on an electronic display. To qualify for this category, the images must contain GUI elements that are unambiguous and recognizable. While a screenshots are a type of complex image, they are broken out here because of their prevalence and the differences in the investigatory approach compared to general complex images.
@@ 5.4 Manipulation Types
  • Text-based images do not require the visual content of an image to possess any specific qualities. They are misinformation claims, as text, rendered in an image. The image may feature other content, but that content is not important for the claim being made. Images which feature no misinformation-relevant non-textual visual content at all are a common element of this category.
@@ 5.4.1 Content Manipulations
Content manipulations are defined here as taking place when (i) the content of the image itself has been altered or synthesized, (ii) it appears endogenous or “native” to the image itself (i.e., not clearly overlaid or added after-the-fact [^15] This is a generous criteria; even very crude attempts to make the manipulations appear present in the original are included here.) and (iii) the alteration made is relevant to the misinformation claim. Content manipulations have existed since the inception of photography, and are perhaps the canonical form of image-based misinformation in popular consciousness. They are created using a very wide variety of techniques, which are loosely situated onto a spectrum between Deepfakes (i.e., AI-generated) and Cheapfakes (“classical,” pre-AI methods) [26].
@@ 5.4.4 Fake Official Documents
This category is characterized by the presence of logos, letterhead, or other attributes that lend the false appearance of an official communication by a government, company, or other well-known organization or official body. They may also appear as articles from reputable news sources, either as physical news clippings or screenshots. Examples of fake official documents are given in Fig. 9. More detail on this category is available in the appendix (Sec. B.3.1).
@@ 6.2 Raters
A total of 83 raters participated in the study. Because of the complexity of the task, careful training and retention was prioritized over a large rater pool. Mean rater tenure was 147 days, performing an average of 5,971 annotations (counting annotations that were subsequently dropped or excluded from the analysis due to changes methodology). Two groups of raters were used simultaneously, one located in India and one in the United States. The groups were preferentially allocated fact checks sourced from their respective countries so that annotations would benefit from contextual knowledge; fact checks from neither country were allocated randomly.
@@ 7.1 Use of Media in Misinformation Claims
Fig. 11 plots the percentage of misinformation claims that rely on media using a 120-day sliding window based on an approximate publication date. While the percentage of media-based misinformation claims among all misinformation claims has been largely stable at least since 2019, such claims represent a large majority of around 80%, at least among those that are subject to a fact check. The cause of the decline between 2017 and 2019 is unclear; and may be due to changing prioritization in content addressed by fact checkers. The date of the fact check is used to reckon the date of the misinformation claim in this plot, and all subsequent plots containing data information. The fact check date is based on a combination of self-reported dates provided by fact checkers as part of the ClaimReview markup schema and the date when the ClaimReview-bearing fact check was encountered [^16] The dates provided in the dataset release are based purely on the self-reported dates from the ClaimReview schema..
@@
The secular increase in video-based misinformation is most apparent as a windowed percentage of total media-based misinformation addressed in fact checks over time. Viewed cumulatively in Fig. 13 the rise in video-based misinformation does not appear to occur alongside a similarly strong deceleration in other modalities. This suggests increases in fact checker capacity is preferentially allocated to video-based misinformation.
@@ 7.3 Manipulation Types
Our study provides no measure of the comparative effectiveness or impact of these methods. Despite this, we do not regard the simplicity of producing a particular manipulation as a measure of its ineffective or benign nature. Rather, the consistent popularity of context manipulation suggests their enduring effectiveness as a means of making a compelling false claim.
@@ 7.4 Content Manipulations
The prevalence of content manipulation sub-types (Sec. 5.4.1), as measured against the proportion of content manipulation types overall, is shown in Fig. 18 and has remained largely stable up until 2023. General (e.g. so-called “photoshops”) and text (applied specifically edits made to text that appears to occur on an object in the scene) manipulations are similarly common and make up a significant majority. Chyron manipulations are a relatively small proportion overall. AI-generated images made up a minute proportion of content manipulations overall until early last year. Starting shortly before 2023, generative AI-generated images begin to rapidly rise as a proportion of overall fact-checked image manipulations; such that the aggregate population of AI-generated images is now far greater than that of manipulated chyrons and nearly as common as text and general content manipulations. Note that the accuracy of the total count of fact-checked AI-content is subject in part to the completeness of their corresponding fact checks (Sec. 9.1): if the fact check does not indicate it’s AI, it will not be recorded as such by the rater.
@@
Figure 19: The “Hype Index” among content manipulations types in image-based misinformation over time. The “Hype Index” is computed as the mean number of fact checks per image over time, separated by content manipulation type. Inset is an AI-generated image of Pope Francis in a white jacket, which went viral in March of 2023 and was subject to numerous independent fact checks as well as being widely featured in the press in general. We interpret this the number of fact checks an image is subject to as an indicator of the fact checker community’s view of the importance of fact checking it. Under this interpretation, the massive spike in the index for AI-generated content indicates that fact checkers viewed communicating the nature of these images was of particular public interest (or at least, of interest to the public) in early 2023. Plots are otherwise computed as in Fig. 11, although the minimum number of datapoints per interval required for inclusion has been changed from 100 to 10 to highlight the sharp spike in fact checks per AI image.
@@ 7.5 Context Manipulations
Images that are asserted by the fact check to be subject to context manipulations but which do not fall into the categories defined in Sec. 5.4.2 are not further characterized, and account for about 40% of all context manipulations. This reflects relatively stringent criteria for inclusion in a subcategory, and the breadth of what is considered context (more complete categorization of context manipulations is identified as an avenue for further research in Sec. 9.5). For example, a context manipulation can falsely state the identity of the originator of an image, but not of anything in the image, and so it would not fall under the identity manipulation subcategory [^17] Not capturing these additional subtypes of context manipulation is due mostly to the need to contain the complexity of the overall rater task.. The prevalence of these categories are shown in Fig. 20.
@@ 7.6 Image Text
Raters noted the presence of text, and did not discriminate between cases where the text occurs on an object in the scene or is digitally overlaid on top of the image. When text was both present and legible, or when a transcript was provided by the fact check (especially in the event the text was in a non-English language), raters were asked to assess whether or not the text was also relevant to the misinformation. Text relevance was determined using criteria similar to those used for the relevance of the image itself: some component of the misinformation claims depends on, or is supported by, the text content in the image. This dependence need not be particularly strong, and the category is used to separate cases where all text that occurs in an image is incidental to the misinformation claim it is associated with.
@@
Figure 22: Self-contextualizing images. The prevalence of ‘self-contextualizing images,’ images subject to context manipulations where the text articulating the false context is present in the image itself, among the total population of context manipulations. Plot created as in Fig. 11.
@@ 7.8 Fake Official Document
Fake official documents appear to be or to contain an official communication of some form from a well-known (or made to appear well-known) and reputable organization. The frequency of fake official documents is plotted in Fig. 24. Such images are of concern as they can disseminate false information during crises, masquerading as an official dispatch containing important information. Particularly salient examples occurred during the COVID-19 pandemic, where misinformation-bearing images would regularly occur appearing to be official information or instruction from a public health body. “Vaccine exemption cards” and the like also fall into this category.
@@ 7.10 Effects of Global Events
The volume of fact checks correlates somewhat with global events. The period between the first recorded instance of COVID-19 in the United States on Jan 20, 2020 and the administration of the first of the vaccines on December 14, 2020 corresponds to the most rapid increase in fact checks since the introduction of ClaimReview. To measure the effect on the distribution of different media types, we checked a battery of notable events against the observed trends. We find little evidence that the events themselves drive shifts in the composition of the misinformation-associated media population. Rather, trends appear to be driven by other phenomena, like the increasing popularity of video platforms or availability of generative-AI methods.
In some cases, changes are observed consistent with a priori expectations. For instance, the “Fake Official Documents” category experiences spikes in prevalence following the deployment of COVID-19 tests and vaccinations. However, these observations are by no means conclusive.